cquation The actu aur Cnot avcilab ut the folloui よata : 22 너 a) what i b.? b) what is bo
Answer #1

(a) We have b_1 = rac{sum (X-ar{X})(Y-ar{Y})}{sum (X-ar{X})^2} or b_1 = rac{sum xy}{sum x^2} , for x be the mean deviation of X values and y be the mean deviation of Y values. Putting the values, we have -654 or bi =-10 .

(b) As the mean is not given, neither is the number of observation, we must find the number of observations n.
We have

sum (X - ar{X})^2 = sum x^2

or sum (X^2 - 2Xar{X} + ar{X}^2) = 60

or sum X^2 - 2ar{X}sum X + ar{X}^2sum 1 = 60

or sum X^2 - 2ar{X}sum X + nar{X}^2 = 60

or sum X^2 - 2(rac{sum X}{n})sum X + n(rac{sum X}{n})^2 = 60

or sum X^2 - rac{2(sum X)^2}{n} + rac{n(sum X)^2}{n^2} = 60

or sum X^2 - rac{2(sum X)^2}{n} + rac{(sum X)^2}{n} = 60

or sum X^2 - rac{(sum X)^2}{n} = 60

or 310 60

or 310 - 60 = rac{(50)^2}{n}

or 2500 250

or n = 10 .

The intercept coefficient would be hence b_0 = ar{Y} - b_1 ar{X} or b_0 = rac{sum Y}{n} - b_1 rac{sum X}{n} or b_0 = rac{580}{10} - 10.9*rac{50}{10} or 0 = 3.5 .